Grade 12Physical Chemistry

The total vapour pressure of 4 mole percent solution of ammonia in water at 293K is 50torr ,the vapour pressure of pure water is 17 torr at this temperature .Apply Henry and RRaults law calculate the total Vapour pressure of 5 mole percent solution.

Kirtan shukla
8 Years agoGrade 12

1 Answer

Sandeep Bhamoo
ApprovedApproved Tutor Answer8 Years ago
Let vapour pressure of pure ammonia=x, then ,50=17(0.96) +x(0.04), x=842torr. Now new mole fraction are 0.95(water), 0.05(ammonia). Hence P= 0.95(17) +0.05(842) =58.25 torr. Please do approve my answer if correct